4. A multi-programmed operating system running on a single CPU assigns time slices of CPU time to the virtual machines t...

4.

A multi-programmed operating system running on a single CPU assigns time slices of CPU time to the virtual machines that run the individual programs. Each virtual machine is given access to the CPU for a certain time (the time slice), during which it will continue to execute its program. The next time slice then goes to another virtual machine, and so on.

a.

Assume that each time slice is one millisecond (1/1000 of a second) long. The computer on which the operating system runs therefore has 1000 time slices available. If the OS itself uses 100 of these (you may assume that these are spread at regular intervals over each second) it can make 900 such time slices available for virtual machines each second. If the computer needs to run 10 programs that each require a time slice at least 20 times per second, and one program (the user interface) that requires a time slice at least once every 10 milliseconds, what is the maximum number of time slices left for other programs? [5 marks]

b.

Given the scenario above, is it theoretically possible for a program that requires 90 time slices to complete in 100 ms or less? Explain. Hint: You will need to consider when exactly the programs under (a) above may need to occupy available time slices. [5 marks]

c.

If the computer had a second CPU, each of the CPUs could process a time slice in parallel and therefore there would be another 1000 time slices per second. Suppose the operating system could make another 900 of these extra time slices available. Would this make it possible for a program that requires 90 time slices to complete in less than 80 ms? Explain. [5 marks]
Answer #1

answer:

a)

1 second has 1000 milliseconds in it.

so this means we have 1000 time slices available

out of these 100 are occupied by the operating system

and 900 are available for application programs

To make a sure the maximum number of time slices remaining we must a lot the least amount of time slices to the given programs

for 10 programs at least 20 times per second will require= 20*10= 200-time slices at least

so now, we have 900-200= 700-time slices available

Now we need a time slice at least once every 10 milliseconds and

there are 100 10 milliseconds in a 1-second time slot.

This means we have to assign 100 more time slice

Now the remaining time slices= 700-100= 600 time slices

b)

Yes, since there are 100 10 ms in a second which means 100 time slices are required here.

c)

No, it won't be possible in this case.

Please give it a thumbs up if this helped you, also provide your valuable feedback.

Please give it a thumbs up if this helped you, also provide your valuable feedback.